OpenStudy (anonymous):

Can this equation be solved? -3|6n-2|+5=8 ill give u a medal plz

OpenStudy (anonymous):

hint -3|6n-2|+5=8 -3|6n-2|+5=8 -3|6n-2|=8 -5 -3|6n-2|=3 so |6n-2| = -1 use |x| = x if x>=0 = -x if x<0 or if |x|=5 then x=5,-5

OpenStudy (anonymous):

your seriously a life saver lmao

OpenStudy (anonymous):

so the finl answer is 5,-5 right?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

|6n-2| = -1 do from here

OpenStudy (anonymous):

i dont know how.. would it be no solution?

OpenStudy (anonymous):

oh would it be n=1/6??

OpenStudy (anonymous):

can u hlp then